solve by completing the square


9x^2 +18X=-5


can anyone give a full break down why the answer is -1/3, -5/3 confused on how my book got -1/3 as an answer.

Answer by stanbon(75887)

solve by completing the square
9x^2 +18X=-5
---
9(x^2+2x+1) = -5+9
---
9(x+1)^2 = 4
---
(x+1)^2 = 4/9
----
x+1 = +2/3 or x+1 = -2/3
---
x = -1/3 or x = -5/3
============================
Cheers,
Stan H.



can anyone give a full break down why the answer is -1/3, -5/3